Question.
1- Market research indicates that you can sell 40 000 tickets for the Sundowns-pirates clash at R10 each, or 30 000 tickets at R20 each .which option would you choose? What is the price elasticity of the demand for tickets for for this particular game? 5 (marks)
1
Expert's answer
2020-03-19T11:50:53-0400

"revenue =price\times quantity"

Options one total revenue is "40000\times 10=400000" , option two total revenue is "30000\times20=600000" . Therefore, I would choose option two due to the high revenue.

"PED=\frac{\Delta{Q}}{\Delta{P}}\times\frac{P}{Q}"

"=\frac{-10000}{10}\times{\frac{10}{40000}}=-0.25" . The negative sign shows the direction of change. Since the elasticity is less than one, the demand for the tickets is price inelastic and one can earn more by increasing the price.